OpenStudy (anonymous):

Solve for x, where x is a real number 3√ (x-2) -8 = 8 Could any1 please explain how to solve this step by step. I really want to be able to understand this. I already know x=274/9

jimthompson5910 (jim_thompson5910):

Is the equation \[\Large \sqrt[3]{x-2}-8 = 8\] OR is the equation \[\Large 3*\sqrt{x-2}-8 = 8\]

OpenStudy (solomonzelman):

(whatever it is) \(\normalsize\color{black}{ \sqrt[3]{x-2}-8=8}\) \(\normalsize\color{black}{ \sqrt[3]{x-2}=0}\) \(\normalsize\color{black}{ \sqrt[3]{x-2}=\sqrt[3]{0}}\) \(\normalsize\color{black}{ x-2=0}\) \(\normalsize\color{black}{3\sqrt{x-2}-8=8}\) \(\normalsize\color{black}{ 3\sqrt{x-2}=0}\) \(\normalsize\color{black}{ \sqrt{x-2}=0}\) \(\normalsize\color{black}{ x-2=0}\)

OpenStudy (solomonzelman):

It doesn't matter JT :)

jimthompson5910 (jim_thompson5910):

true, didn't think that far ahead

OpenStudy (anonymous):

@jim_thompson5910 It's the first one

jimthompson5910 (jim_thompson5910):

actually no

jimthompson5910 (jim_thompson5910):

you have to add 8 to both sides, not subtract 8 from both sides

OpenStudy (solomonzelman):

Ohh, dang yeah... Silly mistake -:(

jimthompson5910 (jim_thompson5910):

Did you get this far andreakristina? \[\Large \sqrt[3]{x-2}-8 = 8\] \[\Large \sqrt[3]{x-2}-8+8 = 8+8\] \[\Large \sqrt[3]{x-2} = 16\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what's the next step?

OpenStudy (anonymous):

@jim_thompson5910 oh wait my bad the 3 is actually multiplying the sqrt

jimthompson5910 (jim_thompson5910):

so we have \[\Large 3*\sqrt{x-2} = 16\]

jimthompson5910 (jim_thompson5910):

what's the next step?

OpenStudy (anonymous):

um, no idea. would we divide 3 from both sides,or do we sqrt the 16?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large 3*\sqrt{x-2} = 16\] \[\Large \sqrt{x-2} = \frac{16}{3}\]

jimthompson5910 (jim_thompson5910):

then we square both sides we square both sides to get rid of the square root

jimthompson5910 (jim_thompson5910):

\[\Large 3*\sqrt{x-2} = 16\] \[\Large \sqrt{x-2} = \frac{16}{3}\] \[\Large (\sqrt{x-2})^2 = (\frac{16}{3})^2\] \[\Large x-2 = \frac{256}{9}\]

jimthompson5910 (jim_thompson5910):

hopefully you see what the last step has to be?

OpenStudy (anonymous):

@jim_thompson5910 I add the two to the fraction correct. thank u so much i completely understand it now :)

jimthompson5910 (jim_thompson5910):

correct, you add 2 to both sides

OpenStudy (anonymous):

@jim_thompson5910 could u help me with one more? 27^(2x)=9^(x-3)